[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri May 27, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri May 27 12:26:58 UTC 2016


HSU Shift Log for Fri May 27 2016 08:26:56 GMT-0400 (EDT)

Shift: H

Forecaster: D Roberts

Technical Issues Summary:

ATCF issues: 1145z TAFB fix did import correctly, but the SAB fix did not.

N-AWIPS/AWIPS2 issues: none

Other Techincal issues: none
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160527/49121a66/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list